A ship with a mass of 52 000 metric tons is capable of producing a thrust force of 25 800 kN while a total resistive force of 5 000 kN is against it, what is the acceleration of the ship? *


1
Expert's answer
2022-01-28T08:14:12-0500

Given:

"m=52\:000\:000\:\rm kg"

"F_{\rm thrust}=25\:800\:000\:\rm N"

"F_{\rm resistive}=5\:000\:000\:\rm N"


The Newton's second law says

"a=\frac{F_{\rm net}}{m}=\frac{F_{\rm thrust}-F_{\rm resistive}}{m}"

"a=\frac{25800000-5000000}{52000000}=0.4\:\rm m/s^2"